The stock of Payout Corp. will go ex-dividend tomorrow. The dividend will be $2.25 per share, and there are 50,000 shares of stock outstanding. The market-value balance sheet for Payout is shown below.


Assets Liabilities and Equity
  Cash $450,000        Equity $1,650,000     
  Fixed assets 1,200,000     
a. What price is Payout stock selling for today? (Round your answer to the nearest cent.)
  Price    $

       

b. What price will it sell for tomorrow? Ignore taxes. (Round your answer to 2 decimal places.)
  Price    $
Answer #1

a: Price of stocks today = Market value of equity/ Number of shares + dividend

= 1650,000/ 50000 + 2.25

= $33+ 2.25

= $ 35.25

b: Price of stocks tomorrow = 35.25 - 2.25 = $33